# Quantum mechanics, non-locality, and the space discreteness hypothesis

### v1

@@ Abstract.
[…]
@@ 1. Introduction
In the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement is greater than or equal to the Planck length, [7]. The interpretation of this inequality is that there are no intervals, just points below the Planck length. We argue that this last assertion can be reformulated in precise mathematical form as: at the Planck length, the physical space can be modeled as a completely disconnected space $\mathcal{X}$. In this framework, there is no continuous world line joining two different points in space $\mathcal{X}$; consequently, the assumption that $\mathcal{X}$ is a model of space is entirely incompatible with the theory of relativity. In this paper,  $\mathcal{X}$ is the mathematical realization of the hypothesis of the discreteness of space at small distances. The naive idea of using a lattice in $\mathbb{R}^{3}$ as a model for the space discreteness is not useful at all.
We propose $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{3}$ as a model of the space-time (or more generally, $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{N}$ as a model for the configuration space). The advantage of this model is that it contains regions homeomorphic to $\mathbb{R}\times\mathbb{R}^{3}$ and to $\mathbb{R}\times\mathcal{X}^{3}$. In the first type of regions, causal and local theories like the relativistic QM and TR can be formulated. In contrast, in the second type of regions, acausal and non-local theories can be formulated. In this framework,  QM (in sense of Dirac-von Neumann) on the Hilbert spaces $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ can be rigorously formulated, [8]-[11]. A rigorous mathematical study of QM on Hilbert spaces of type $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ is possible assuming suitable technical conditions, but this choice is quite involved. Here, we take $N=1$, $\mathbb{Q}_{p}$, the field of $p$-adic numbers; $\mathbb{Q}_{p}$ is a paramount example of a completely disconnected space, with a very rich mathematical structure. In the following paragraphs, we provide an argument showing the physical relevance of the chosen model.
@@ 2.2. Macroscopic versus microscopic
We divide the space-time into two realms: the macroscopic and the microscopic. In the first one, the model of the space-time is $\mathbb{R}\times\mathbb{R}^{3}$, while in the second one, it is $\mathbb{R}\times\mathcal{X}^{3}$, where $\mathcal{X}$ is a completely disconnected topological space. The connectedness of a topological space is a topological property, which means it remains invariant under continuous transformations (homeomorphisms). Therefore, $\mathcal{X}$ is not topologically equivalent to $\mathbb{R}$, nor a proper subset of $\mathbb{R}$.
@@ 2.3. The $p$-adic space
In this section, we fix the notation and collect some basic results on $p$-adic analysis that we use in this paper. For a detailed exposition on $p$-adic analysis, the reader may consult [15], [40], [41], [42]. For a discussion on applications of $p$-adic analysis to physics, biology, computing, and other areas, the reader may consult [19], [22], [25], [43], [44], [45], [46], [48], and the references therein.
@@
with $x_{-k}\neq 0$, where $k$ is an integer, and the $x_{j}$s are numbers from the set $\left\{0,1,\ldots,p-1\right\}$. The set of all possible sequences of the form (2.2) constitutes the field of $p$-adic numbers $\mathbb{Q}_{p}$. There are natural field operations, sum and multiplication, on series of form (2.2). There is also a norm in $\mathbb{Q}_{p}$ defined as $\left|x\right|_{p}=p^{-ord(x)}$, where $ord_{p}(x)=ord(x)=k$, for a nonzero $p$-adic number $x$. By definition $ord(0)=\infty$. The field of $p$-adic numbers with the distance induced by $\left|\cdot\right|_{p}$ is a complete ultrametric space. The ultrametric property refers to the fact that $\left|x-y\right|_{p}\leq\max\left\{\left|x-z\right|_{p},\left|z-y\right|_{p}\right\}$ for any $x$, $y$, $z$ in $\mathbb{Q}_{p}$. The $p$-adic integers which are sequences of form (2.2) with $k\geq 0$. All these sequences constitute the unit ball $\mathbb{Z}_{p}$. The unit ball is an infinite rooted tree with fractal structure. As a topological space $\mathbb{Q}_{p}$ is homeomorphic to a Cantor-like subset of the real line, see, e.g., [15], [42]. $\mathbb{Q}_{p}$ is a paramount example of a completely disconnected space, with a very rich mathematical structure.
@@ 2.5. Non-locality in $p$-adic QM
Along this paper, we use discrete space as a synonymous of completely disconnected space. The non-locality is natural in discrete spaces like $\mathbb{Q}_{p}$. On a completely disconnected topological space, like $\mathbb{Q}_{p}$, the locally constant functions play a central role. A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is locally constant, if $\varphi\left(x+x^{\prime}\right)-\varphi\left(x\right)=0$ for $x^{\prime}$ sufficiently small. Then, the standard notion of derivative defined as
@@ 3. $p$-Adic QM is the quantum model for quantum computing
We denote by $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ the $N$-dimensional Hilbert space spanned by (3.1). Then, $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ is isomorphic to $\mathbb{C}^{N}$, as Hilbert spaces. Notice that the isomorphism exists for any $p$ and $l$ such that $N\leq p^{l}$.
The advantage of using $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ over $\mathbb{C}^{N}$ is that we can construct, in a natural way, continuos versions of quantum models in $\mathbb{C}^{N}$. We identify $p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ with $e_{I}$, where $\left\{e_{I}\right\}_{I\in G_{l}^{0}}$ is the canonical basis of $\mathbb{C}^{N}$. Let
@@
is a continuous Schrödinger equation attached to matrix $\left[H_{J,K}\right]_{1\leq J,K\leq N}$. For further details, the reader may consult [21]; see also [24], [25], and the references therein.
@@ 4.2. Some results about tensor product of Hilbert spaces
We set $L^{2}(\mathbb{R},dx_{\infty})=L^{2}(\mathbb{R})$, where  denotes the Lebesgue measure, and $L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)=L^{2}\left(\mathbb{Q}_{p}\right)$, where $dx_{p}$ denotes the Haar measure of $\mathbb{Q}_{p}$. We denote by $dx_{\infty}dx_{p}$, the product measure of $dx_{\infty}$ and $dx_{p}$ on $\mathbb{R}\times\mathbb{Q}_{p}$. On the other hand, $L^{2}\left(\mathbb{Q}_{p}\right)$ is a separable Hilbert space because it has a countable orthonormal basis $\left\{\psi_{rbk}\left(x_{p}\right)\right\}_{rbk}$, and also $L^{2}(\mathbb{R})$ is a separable Hilbert space. The Haar wavelets $\left\{H_{m}(x_{\infty})\right\}_{m\in\mathbb{N}}$, which are functions with compact support, form an orthonormal basis of $L^{2}(\mathbb{R})$; also the Hermite polynomials, which are Schwartz functions in $\mathbb{R}$, form an orthonormal basis of $L^{2}(\mathbb{R})$. Then $\left\{H_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right)\right\}_{rbk,m}$ is an orthonormal basis of $L^{2}(\mathbb{R})\otimes L^{2}\left(\mathbb{Q}_{p}\right)$, cf. [54, Proposition 2.7] or [55, Section II.4, Proposition 2], where $\otimes$ denotes the tensor product of Hilbert spaces.
@@ 4.4. Non-local Schrödinger equations
$\mathbb{R}\times\left(\left\{\beta\right\}\times\mathbb{Q}_{p}\right)^{N}\simeq\mathbb{R}\times\mathbb{Q}_{p}^{3N},$
@@ 5. The measurement problem and GWR theory
In this section, we review quickly the measurement problem and the Ghirardi-Rimini-Weber (GRW) collapse theory. In the next section, we propose a solution to the measurement problem; the collapse mechanism resembles the one in the GRW theory, but there are several fundamental differences.
@@ 5.2. The GRW collapse model
The parameter $r$ is random with a probability distribution $P(r)$. The localization (5.2) occurs at random times, with a Poisson distribution with mean frequency $\lambda$. The parameters $\lambda$ and $\sigma$ are phenomenological constants that should be understood as new constants of Nature. The theory can be formulated through a short list of principles, from these using the statistical operator formalism, a master equation for $N$ particles can be obtained. This equation is a substitute to the standard Schrödinger equation; see [35], and the references therein.
@@ 8. Conclusions and final discussion
QM on $\mathbb{C}^{N}$ is a particular case of $p$-adic QM, more precisely, it is QM on $L^{2}(\chi_{N}\left(\mathbb{Z}_{p}\right))$. $p$-Adic QM is non-local because the Hamiltonians (non-local operators) allow spooky action at a distance. Thus, this theory admits realism. The implementation of the space discreteness hypothesis requires using a completely disconnected space $\mathcal{X}$ as a model of the physical space at the microscopic level. In such a space, there are no continuous curves (world lines), so any theory using this type of space is not compatible with the theory of relativity. It does not imply any immediate prediction on the violation of Einstein’s causality.

### v2

@@ Abstract.
The space discreteness hypothesis asserts that the nature of space at short distances is radically different from that at large distances. Based on the Bronstein inequality, here, we use a totally disconnected topological space $\mathcal{X}$ as a model for the physical space. However, we consider the time as a real variable. In this framework, the Dirac-von Neumann formalism can be used. This discreteness hypothesis implies that given two different points in space, there is no continuous curve (a world line) joining them. Consequently, this hypothesis is not compatible with the theory of relativity. We propose $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{3}$ as a model of a space-time. For simplicity, we work out our models with $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)$ as the configuration space. Quantum mechanics (QM), in the sense of Dirac-von Neumann, on the Hilbert space $L^{2}\left(\mathbb{R}\times\mathcal{X}\right)$ is a non-local theory: the Hamiltonians are non-local operators, and thus, spooky action at a distance is allowed. The paradigm asserting that the universe is non-locally real implies that the proposed version of QM admits realism. This version of QM can be specialized to standard QM by using Hamiltonians acting on wavefunctions supported on the region $\mathbb$$\times R$}. We apply the developed formalism to the measurement problem. We propose a new mechanism for wavefunction collapse. The mechanism resembles that proposed by Ghirardi, Rimini, and Weber, but there are significant differences. The most important feature is that the Schrödinger equation describes the dynamics at all times, even at the moment of measurement. We also discuss a model for the two-slit experiment in which bright and dark states of light (recently proposed) naturally occur.
@@ 1. Introduction
In the 1930s, Bronstein showed that general relativity and quantum mechanics imply that the uncertainty $\Delta x$ of any length measurement is greater than or equal to the Planck length, [7]. A widely accepted interpretation of this inequality is that there are no intervals, just points below the Planck length. We argue that this last assertion can be reformulated in precise mathematical terms as follows: at the Planck length, the physical space can be modeled as a totally disconnected space $\mathcal{X}$. In this framework, there is no continuous world line joining two different points in space $\mathcal{X}$; consequently, the assumption that $\mathcal{X}$ is a model of space is entirely incompatible with the theory of relativity. In this paper, $\mathcal{X}$ is the mathematical realization of the hypothesis of the discreteness of space at small distances. The naive idea of using a lattice in $\mathbb{R}^{3}$ as a model for the space discreteness is not useful at all.
We propose $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{3}$ as a model of the space-time (or more generally, $\mathbb{R}\times\left(\mathbb{R}\times\mathcal{X}\right)^{N}$ as a model for the configuration space). The advantage of this model is that it contains regions homeomorphic to $\mathbb{R}\times\mathbb{R}^{3}$ and to $\mathbb{R}\times\mathcal{X}^{3}$. In the first type of regions, causal and local theories like the relativistic QM and TR can be formulated. In contrast, in the second type of regions, acausal and non-local theories can be formulated. In this framework, QM (in sense of Dirac-von Neumann) on the Hilbert spaces $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ can be rigorously formulated, [8]-[11]. A rigorous mathematical study of QM on Hilbert spaces of type $L^{2}\left(\left(\mathbb{R}\times\mathcal{X}\right)^{N}\right)$ is possible assuming suitable technical conditions, but this choice is quite involved. Here, we take $N=1$, $\mathcal{X}=\mathbb{Q}_{p}$, the field of $p$-adic numbers; $\mathbb{Q}_{p}$ is a paramount example of a totally disconnected space, with a very rich mathematical structure. In the following paragraphs, we provide an argument showing the physical relevance of the chosen model.
@@ 2.2. Macroscopic versus microscopic
We divide the space-time into two realms: the macroscopic and the microscopic. In the first one, the model of the space-time is $\mathbb{R}\times\mathbb{R}^{3}$, while in the second one, it is $\mathbb{R}\times\mathcal{X}^{3}$, where $\mathcal{X}$ is a totally disconnected topological space. The connectedness of a topological space is a topological property, which means it remains invariant under continuous transformations (homeomorphisms). Therefore, $\mathcal{X}$ is not topologically equivalent to $\mathbb{R}$, nor a proper subset of $\mathbb{R}$.
@@ 2.3. The $p$-adic space
In this section, we fix the notation and collect some basic results on $p$-adic analysis that we use in this paper. For a detailed exposition on $p$-adic analysis, the reader may consult [15], [40]-[42]. For a discussion on applications of $p$-adic analysis to physics, biology, computing, and other areas, the reader may consult [19], [22], [25], [43]-[46], [48], and the references therein.
@@
with $x_{-k}\neq 0$, where $k$ is an integer, and the $x_{j}$s are numbers from the set $\left\{0,1,\ldots,p-1\right\}$. The set of all possible sequences of the form (2.2) constitutes the field of $p$-adic numbers $\mathbb{Q}_{p}$. There are natural field operations, sum and multiplication, on series of form (2.2). There is also a norm in $\mathbb{Q}_{p}$ defined as $\left|x\right|_{p}=p^{-ord(x)}$, where $ord_{p}(x)=ord(x)=k$, for a nonzero $p$-adic number $x$. By definition $ord(0)=\infty$. The field of $p$-adic numbers with the distance induced by $\left|\cdot\right|_{p}$ is a complete ultrametric space. The ultrametric property refers to the fact that $\left|x-y\right|_{p}\leq\max\left\{\left|x-z\right|_{p},\left|z-y\right|_{p}\right\}$ for any $x$, $y$, $z$ in $\mathbb{Q}_{p}$. The $p$-adic integers which are sequences of form (2.2) with $k\geq 0$. All these sequences constitute the unit ball $\mathbb{Z}_{p}$. The unit ball is an infinite rooted tree with fractal structure. As a topological space $\mathbb{Q}_{p}$ is homeomorphic to a Cantor-like subset of the real line, see, e.g., [15], [42]. $\mathbb{Q}_{p}$ is a paramount example of a totally disconnected space, with a very rich mathematical structure.
@@ 2.5. Non-locality in $p$-adic QM
Along this paper, we use discrete space as a synonymous of totally disconnected space. The non-locality is natural in discrete spaces like $\mathbb{Q}_{p}$. On a totally disconnected topological space, like $\mathbb{Q}_{p}$, the locally constant functions play a central role. A function $\varphi:\mathbb{Q}_{p}\rightarrow\mathbb{C}$ is locally constant, if $\varphi\left(x+x^{\prime}\right)-\varphi\left(x\right)=0$ for $x^{\prime}$ sufficiently small. Then, the standard notion of derivative defined as
@@ 3. $p$-Adic QM is the quantum model for quantum computing
We denote by $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ the $N$-dimensional Hilbert space spanned by (3.1). Then, $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ and $\mathbb{C}^{N}$ are isomorphic as Hilbert spaces. Notice that the isomorphism exists for any $p$ and $l$ such that $N\leq p^{l}$.
The advantage of using $\mathcal{\chi}_{N}(\mathbb{Z}_{p})$ over $\mathbb{C}^{N}$ is that we can construct, in a natural way, continuous versions of quantum models in $\mathbb{C}^{N}$. We identify $p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-I\right|_{p}\right)$ with $e_{I}$, where $\left\{e_{I}\right\}_{I\in G_{l}^{0}}$ is the canonical basis of $\mathbb{C}^{N}$. Let
@@
is a continuous Schrödinger equation attached to matrix $\left[H_{J,K}\right]_{1\leq J,K\leq N}$. For further details, the reader may consult [21]; see also [24]-[25], and the references therein.
@@ 4.2. Some results about tensor product of Hilbert spaces
We set $L^{2}(\mathbb{R},dx_{\infty})=L^{2}(\mathbb{R})$, where $dx_{\infty}$ denotes the Lebesgue measure, and $L^{2}\left(\mathbb{Q}_{p},dx_{p}\right)=L^{2}\left(\mathbb{Q}_{p}\right)$, where $dx_{p}$ denotes the Haar measure of $\mathbb{Q}_{p}$. We denote by $dx_{\infty}dx_{p}$, the product measure of $dx_{\infty}$ and $dx_{p}$ on $\mathbb{R}\times\mathbb{Q}_{p}$. On the other hand, $L^{2}\left(\mathbb{Q}_{p}\right)$ is a separable Hilbert space because it has a countable orthonormal basis $\left\{\psi_{rbk}\left(x_{p}\right)\right\}_{rbk}$, and also $L^{2}(\mathbb{R})$ is a separable Hilbert space. The Haar wavelets $\left\{H_{m}(x_{\infty})\right\}_{m\in\mathbb{N}}$, which are functions with compact support, form an orthonormal basis of $L^{2}(\mathbb{R})$; also the Hermite polynomials, which are Schwartz functions in $\mathbb{R}$, form an orthonormal basis of $L^{2}(\mathbb{R})$. Then $\left\{H_{m}\left(x_{\infty}\right)\psi_{rbk}\left(x_{p}\right)\right\}_{rbk,m}$ is an orthonormal basis of $L^{2}(\mathbb{R})\otimes L^{2}\left(\mathbb{Q}_{p}\right)$, cf. [54, Proposition 2.7] or [55, Section II.4, Proposition 2], where $\otimes$ denotes the tensor product of Hilbert spaces.
@@ 4.4. Non-local Schrödinger equations
$\mathbb{R}\times\left(\left\{\beta\right\}\times\mathbb{Q}_{p}\right)^{N}\simeq\mathbb{R}\times\mathbb{Q}_{p}^{N},$
@@ 5. The measurement problem and GWR theory
In this section, we review quickly the measurement problem and the Ghirardi-Rimini-Weber (GRW) collapse theory. In the next section, we propose a solution to the measurement problem. The new collapse mechanism resembles the one in the GRW theory, but there are several fundamental differences.
@@ 5.2. The GRW collapse model
The parameter $r$ is a random variable with a probability distribution $P(r)$. The localization (5.2) occurs at random times, with a Poisson distribution with mean frequency $\lambda$. The parameters $\lambda$ and $\sigma$ are phenomenological constants that should be understood as new constants of Nature. The theory can be formulated through a short list of principles, from these using the statistical operator formalism, a master equation for $N$ particles can be obtained. This equation is a substitute to the standard Schrödinger equation; see [35], and the references therein.
@@ 8. Conclusions and final discussion
QM on $\mathbb{C}^{N}$ is a particular case of $p$-adic QM, more precisely, it is QM on $L^{2}(\chi_{N}\left(\mathbb{Z}_{p}\right))$. $p$-Adic QM is non-local because the Hamiltonians (non-local operators) allow spooky action at a distance. Thus, this theory admits realism. The implementation of the space discreteness hypothesis requires using a totally disconnected space $\mathcal{X}$ as a model of the physical space at the microscopic level. In such a space, there are no continuous curves (world lines), so any theory using this type of space is not compatible with the theory of relativity. It does not imply any immediate prediction on the violation of Einstein’s causality.
@@
Finally, we mention that in adelic QM, the configuration space is the product $\mathbb{R}\times\mathbb{A}_{\mathbb{Q}}$, where $\mathbb{A}_{\mathbb{Q}}$ is the ring of finite adeles, which consists of sequences $\left\{x_{p}\right\}_{p\geq 2}$ from ${\textstyle\prod\nolimits_{p\geq 2}}\mathbb{Q}_{p}$ such that $x_{p}\in\mathbb{Z}_{p}$ for $p$ sufficiently large. Dragovich and his collaborators have studied several quantum models on the infinite-dimensional space $\mathbb{R}\times\mathbb{A}_{\mathbb{Q}}$; see, e.g., [69]-[71]. Extra dimensions in physical models represent a nontrivial feature. Our solution to the measurement problem relies on extra dimensions. As we already noted, the $p$-adic QM describes continuous-time quantum walks, including those that stem from graphs. The physical implementation of these random walks generally involves particles or excitations on networks of size scales ranging from a few nanometers to micrometers. Therefore, it naturally leads the author to ask whether extra dimensions exist at the nanoscale. This fact matches Vafa’s proposal for the existence of at least one extra spatial dimension at the nanoscale, [72].
Data availability statement
No new data were created or analyzed in this study. Data sharing is not applicable to this paper.